A certain reaction obeys second-order kinetics with respect to its only reactant (call it R) and...

  1. A certain reaction obeys second-order kinetics with respect to its only reactant (call it R) and has a rate constant of 0.0520 L mole-1 s-1.  If the initial concentration of reactant R is 0.265 moles/liter, what is the concentration of R after 3.50 minutes?                                                                                                             

_______________
Answer #1

Answer:

The given reaction is

              2R Products

The rate law is -dR/dt = k [R]2

The integrated form of the above rate law is

[ 1/Rt - 1/R0] = kt

Where R0 is the initial concentration of R

          Rt is the concentration of R at time t

Given R0 = 0.265 mol L-1

         t = 3.5 min = 3.5x 60 = 210 s

         k = 0.0520 L mol-1s-1

Therefore 1/Rt = 1/R0 + kt

                     = [1 / 0.265 mol L-1 ]+ 0.0520 L mol-1s-1 x 210 s

                     = 3.774 L mol-1 + 10.92 L mol-1

                     = 14.694 L mol-1

              Rt = [1 / 14.694] mol L-1

                  = 0.068 mol L-1

Thus the concentration of R after 3.5 minutes is 0.068 mol L-1